OpenStudy (anonymous):

given x= 2t^4 + 100^3 ,find dx/dt and find values of t when dx/dt =0

OpenStudy (lgbasallote):

use power rule to find dx/dt

OpenStudy (anonymous):

\[\large \frac{d}{dx}(x)^n = n \cdot (x)^{n-1}\]

OpenStudy (lgbasallote):

in this case.....t

OpenStudy (anonymous):

Put x = t...

OpenStudy (anonymous):

But how do i find x

OpenStudy (lgbasallote):

what do you mean find x?

OpenStudy (anonymous):

you are going to find dx/dt not x..

OpenStudy (anonymous):

I mean to say that first find dx/dt...

OpenStudy (lgbasallote):

here's an example.. \[y = 3x^2 + 4x + 2\] i want to find dy/dx \[\frac{dy}{dx} = 6x + 4\] now i want to find the value/s of x when dy/dx = 0 \[\implies 6x + 4 = 0\] \[\implies 6x = -4\] \[\implies x = -\frac 23\] does that help?

OpenStudy (anonymous):

Which would be 8t^3 +300t^2

OpenStudy (lgbasallote):

\[\frac{dx}{dt} = 8t^3 + 300t^2\] right..now solve for t when dx/dt = 0

OpenStudy (unklerhaukus):

\[\frac{\text d}{\text dt}(t)^n = n \cdot (t)^{n-1}\] \[\frac{\text d}{\text dt}(2t^4 + 100t^3) = \frac{\text d}{\text dt}(2t^4)+ \frac{\text d}{\text dt}(100t^3)\]

OpenStudy (anonymous):

Are you missing t there in your question ??

OpenStudy (lgbasallote):

hmm? i thought you saw that coming?

OpenStudy (anonymous):

Yes you are correct...

OpenStudy (unklerhaukus):

8t^3 +300t^2 \(\checkmark\)

OpenStudy (anonymous):

@UnkleRhaukus \(\large \checkmark\)

OpenStudy (lgbasallote):

@lgbasallote \(\sqrt{}\)

OpenStudy (anonymous):

Ha ha ha ha...

OpenStudy (anonymous):

Now equate what you have found to 0 and try to find values for t..

OpenStudy (anonymous):

But how do i use dx/dt =0 and find value of t

OpenStudy (unklerhaukus):

\[\frac{\text dx}{\text dt}= 8t^3 +300t^2\] \[0= 8t^3 +300t^2\]

OpenStudy (lgbasallote):

hint: you're going to factor something out

OpenStudy (anonymous):

And I don't think its \(t^4\).. Ha ha ha..

OpenStudy (anonymous):

Oh i get it thnx guys

OpenStudy (lgbasallote):

aww i was ready for another example-slash-demo

OpenStudy (unklerhaukus):

what are the values for \(t\) ?

OpenStudy (anonymous):

Dw i llbe posting alot of question which will make you scrach your head

OpenStudy (lgbasallote):

i like scratches

OpenStudy (unklerhaukus):

scratch and sniff?

OpenStudy (anonymous):

And I like your spirit @lgbasallote